OpenStudy (anonymous):

Unit Rate: Drive 168 mi in 2 h 30 min

OpenStudy (anonymous):

\[\frac{168}{2+\frac{30}{60}}\text{=}\frac{336}{5}\text{= }67.2 \text{ mph}\]

OpenStudy (anonymous):

Thanks Robtobey!

OpenStudy (anonymous):

You are welcome.